# Spectral deviation of concentration operators on reproducing kernel Hilbert spaces

### 11 Mar 2026 v1

@@ 8.1.1. Gelfand-Shilov window classes
for constants $B,\alpha>0$ — see [19, 10], [9, Theorem 3.2] or [22, Corollary 3.11 and Proposition 3.12], which gives the kernel estimate $|K(z,w)|=|V_{g}g(z-w)|\leq Be^{-a|z-w|^{1/\beta}}$.
@@
where $\delta\in(0,1)$ and $\tau\coloneqq\max\{\frac{1}{\delta},\frac{1}{1-\delta}\}$. This essentially recovers [39, Theorem 1.1]. (The cited result provides a slightly sharper dependence on $\beta$ when $\Omega$ is subject to increasing dilations. To maintain the conciseness of the present paper, we omit a more detailed discussion of this refinement.)
@@ 8.1.2. Polynomial decay
We now use Theorem 2.2 together with Lemma 9.1 to recover [39, Theorem 1.4] (with a slightly more practical formulation in terms of modulation norms).
@@ 8.2. Gabor multipliers
One of the most important results in Gabor theory is that if $g\in M^{1}(\mathbb{R}^{d})$, then ${g^{\mathrm{d}}_{\Lambda}}\in M^{1}(\mathbb{R}^{d})$ [21, Section 4.2], which intuitively means that the coefficients in the expansion (8.6) reflect the time-frequency profile of the function $f$. See [20, Chaper 5] for more background on Gabor frames.
@@
First let us mention that the frame condition is not a real imposition since it is satisfied for sufficiently small $\sigma$ by Janssen’s criterion. Regarding (i), apply Theorem 7.1 (i) and Remark 8.2 with parameter $2s\geq 1$. It suffices to estimate the quantity $N(2s)$. We compute the modulation space norm (8.4) using the $L^{2}$-normalized Gaussian window $\varphi(x)=2^{d/4}e^{-\pi|x|^{2}}$ and for $1\leq p<\infty$, $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$ we also write
@@
The proof (ii) is quite similar, so we comment on the overall argument and skip the details. We can work directly with $g$ rather than $\varphi$ and proceed as in (8.8) to get
@@
The results follows from Theorem 8.3 and the fact that in dimension 2 connectedness implies Ahlfors regularity with parameters $1\leq\kappa\leq 2$ and $\eta=\mathcal{H}(\partial\Omega)$ (e.g. [40, Lemma 2.5]). ∎
@@ 9.1. Mixed-state localization operators
For $m\in L^{1}(\mathbb{R}^{2d})$ and $R\in\mathcal{S}_{1}$, define their function-operator convolution $m\star R\in\mathcal{S}_{1}$ from quantum harmonic analysis via
@@
Note that Theorem 2.2 combined with the parameters as above improves the estimates of [38, Lemma 4.3] for $\text{trace}\big(A_{\Omega}^{S}-(A_{\Omega}^{S})^{2}\big)$ that correspond to (2.6) when $s=1$ (or more precisely, to (6.2) for $p=2$ and $\alpha=1$). On the other hand, if $S=g\otimes h$, then $A_{\Omega}^{S}=A_{\Omega}^{g}$ from (1.4). So, Lemma 9.1 applies to the scalar case as mentioned in Section 8.1.
@@ 9.2. Fourier concentration operators
The operator $T_{I,J}$ was originally studied in the seminal papers [48, 35, 36] and has proved its usefulness for a variety of applications (see [7, 51, 54, 14]). One-parameter spectral asymptotics for $T_{I,J}$ go back to [34, 37] — see also [53, 49] — while the study of two-parameter spectral asymptotics is much more recent [30, 45, 28, 7]. The higher dimensional case, with general domains in lieu of intervals was treated in [29, 41, 26, 27].
@@
Step 3 (Reproducing kernel estimates). It is easy to check that $\|\widetilde{T}\|\leq 6\eta$ and so,
@@
$1/2\leq\|(K_{n})_{x}\|_{2}^{2}\leq 9\,,\qquad\text{and,}\qquad|K_{n}(x,y)|\leq C_{\beta}^{\prime}e^{-c_{\beta}^{\prime}|\delta_{n}(x-y)|^{1/\beta}}\,$,
where $c_{\beta}^{\prime},C_{\beta}^{\prime}$ do not depend on $\delta_{n}$. Notice that one can renormalize the measure to satisfy [C1], and [C2] clearly holds with $\gamma=1$ and $\operatorname{\nabla}(\Omega)=4$. So, applying Corollary 2.3 to $\Pi_{n}1_{\delta_{n}J}\Pi_{n}$ (and slightly increasing $\beta$ to avoid $\log\log$ terms), we get for a constant $D_{\beta}>0$ independent of $n$,
@@ Acknowledgements
This research was funded by the Austrian Science Fund (FWF) 10.55776/Y1199 (J.L.R., M.S. and L.V.) and 10.55776/PAT1384824 (F.M. and M.S.). F.M. was also supported by the EPSRC UKRI/EP/C003286/1. L. V. acknowledges support from the Erasmus+ Scholarship for Traineeship, issued by the Università di Genova; number 2023/93. This work elaborates in part on the masters’ thesis [52]. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.

### 08 Apr 2026 v2

@@ 8.1.1. Gelfand-Shilov window classes
for constants $B,\alpha>0$ — see [19, 10], [9, Theorem 3.2] or [22, Corollary 3.11 and Proposition 3.12], which gives the kernel estimate $|K(z,w)|=|V_{g}g(z-w)|\leq Be^{-\alpha|z-w|^{1/\beta}}$.
@@
where $\delta\in(0,1)$ and $\tau\coloneqq\max\{\frac{1}{\delta},\frac{1}{1-\delta}\}$. This essentially recovers [40, Theorem 1.1]. (The cited result provides a slightly sharper dependence on $\beta$ when $\Omega$ is subject to increasing dilations. To maintain the conciseness of the present paper, we omit a more detailed discussion of this refinement.)
@@ 8.1.2. Polynomial decay
We now use Theorem 2.2 together with Lemma 9.1 to recover [40, Theorem 1.4] (with a slightly more practical formulation in terms of modulation norms).
@@ 8.2. Gabor multipliers
One of the most important results in Gabor theory is that if $g\in M^{1}(\mathbb{R}^{d})$, then ${g^{\mathrm{d}}_{\Lambda}}\in M^{1}(\mathbb{R}^{d})$ [21, Section 4.2], which intuitively means that the coefficients in the expansion (8.6) reflect the time-frequency profile of the function $f$. See [20, Chapter 5] for more background on Gabor frames.
@@
First, let us mention that the frame condition is not a real imposition since it is satisfied for sufficiently small $\sigma$ by Janssen’s criterion. Regarding (i), apply Theorem 7.1 (i) and Remark 8.2 with parameter $2s\geq 1$. It suffices to estimate the quantity $N(2s)$. We compute the modulation space norm (8.4) using the $L^{2}$-normalized Gaussian window $\varphi(x)=2^{d/4}e^{-\pi|x|^{2}}$ and for $1\leq p<\infty$, $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$ we also write
@@
The proof of (ii) is quite similar, so we comment on the overall argument and skip the details. We can work directly with $g$ rather than $\varphi$ and proceed as in (8.8) to get
@@
The results follows from Theorem 8.3 and the fact that in dimension 2 connectedness implies Ahlfors regularity with parameters $1\leq\kappa\leq 2$ and $\eta=\mathcal{H}(\partial\Omega)$ (e.g. [41, Lemma 2.5]). ∎
@@ 9.1. Mixed-state localization operators
For $m\in L^{1}(\mathbb{R}^{2d})$ and $R\in\mathcal{S}_{1}$, define their function-operator convolution $m\star R\in\mathcal{S}_{1}$ via
@@
Note that Theorem 2.2 combined with the parameters as above greatly improves the estimates of [39, Lemma 4.3] for $\text{trace}\big(A_{\Omega}^{S}-(A_{\Omega}^{S})^{2}\big)$ that correspond to (2.6) when $s=1$ (or more precisely, to (6.2) for $p=2$ and $\alpha=1$). On the other hand, if $S=g\otimes h$, then $A_{\Omega}^{S}=A_{\Omega}^{g}$ from (1.4). So, Lemma 9.1 applies to the scalar case as mentioned in Section 8.1.
@@ 9.2. Fourier concentration operators
The operator $T_{I,J}$ was originally studied in the seminal papers [49, 36, 37] and has proved its usefulness for a variety of applications (see [7, 52, 55, 14]). One-parameter spectral asymptotics for $T_{I,J}$ go back to [35, 38] — see also [54, 50] — while the study of two-parameter spectral asymptotics is much more recent [30, 46, 28, 7]. The higher dimensional case, with general domains in lieu of intervals was treated in [29, 42, 26, 27, 33].
@@
Step 3 (Reproducing kernel estimates). It is easy to check that $\|\widetilde{T}\|\leq 6\eta$ and so
@@
$1/2\leq\|(K_{n})_{x}\|_{2}^{2}\leq 9\,,\qquad\text{and}\qquad|K_{n}(x,y)|\leq C_{\beta}^{\prime}e^{-c_{\beta}^{\prime}|\delta_{n}(x-y)|^{1/\beta}}\,$,
where $c_{\beta}^{\prime},C_{\beta}^{\prime}$ do not depend on $\delta_{n}$. Notice that one can renormalize the measure to satisfy [C1], and [C2] clearly holds with $\gamma=1$ and $\operatorname{\nabla}(\Omega)=4$. Applying Corollary 2.3 to $\Pi_{n}1_{\delta_{n}J}\Pi_{n}$ (and slightly increasing $\beta$ to avoid $\log\log$ terms), we get for a constant $D_{\beta}>0$ independent of $n$,
@@ Acknowledgements
This research was funded by the Austrian Science Fund (FWF) 10.55776/Y1199 (J.L.R., M.S. and L.V.) and 10.55776/PAT1384824 (F.M. and M.S.). F.M. was also supported by the EPSRC UKRI/EP/C003286/1. L. V. acknowledges support from the Erasmus+ Scholarship for Traineeship, issued by the Università di Genova; number 2023/93. This work elaborates in part on the master’s thesis [53]. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission.